# Compression fronts from fast radio bursts

### v1

@@ 6.3 Evolution of a radiative compression front
In the regime of $q\gg 1$, the fluid element moving through the wave packet experiences strong radiative losses. The radio wave sustains the fluid enthalpy $w\approx(3/2)a$ by continually depositing energy to offset the losses (B22). Thus, the received power per particle (measured in the local fluid frame) equals the emitted power ˙$\dot{{\cal E}}_{e}$ (Lorentz invariant). The wave-fluid interaction is basically a scattering process with the cross section $\sigma_{\rm sc}=\dot{{\cal E}}_{e}/F$, where $F=cU$ is the energy flux of the radio wave.
The scattering is accompanied by a radiative pressure force $\dot{\cal P}$, which is Lorentz-invariant and easiest to find in the local fluid frame where $\dot{\cal P}=\dot{{\cal E}}_{e}/c$, so
$\dot{\cal P}=\frac{\dot{{\cal E}}_{e}}{c}=\sigma_{\rm T}\tilde{U}\,\tilde{\gamma}^{2}=\frac{\sigma_{\rm T}U}{\kappa^{2}}\,\tilde{\gamma}^{2}.$ (78)
@@
with $a\sim a_{\max}$. This sets the budget for the fluid momentum exchange with the wave per crossing time $t_{\rm cross}$. The gained momentum is shared with the magnetic field frozen in the fluid, increasing ${\cal P}_{\rm MHD}$.
@@
where $q$ is evaluated using the entire column energy of the wave packet $\Sigma(T)$ (Equation 59). Equation (80) gives the timescale for approaching the steady state, $t_{\star}=t_{\cal P}(\kappa_{\star})$ if we substitute $\kappa\sim\kappa_{\star}\approx qw$ (Equation 58) and $h_{\star}=1+w+\kappa_{\star}\sigma_{\rm u}\approx\kappa_{\star}\sigma_{\rm u}$:
$t_{\star}\sim q\,\sigma_{\rm u}t_{\rm cross}\qquad(q\gg 1,\;\,\sigma_{\rm u}\gg 1).$ (81)
@@
Front relaxation to the steady state in the radiative regime is found by solving Equations (50) and (51) similarly to section 6.2, but now keeping the term $\dot{\gamma}_{e}\neq 0$. This term may be expressed as follows
$\dot{\gamma}_{e}={\cal R}\frac{w^{4}}{\kappa^{2}}\quad(w\gg 1),\qquad{\cal R}\equiv\left(\frac{2}{3}\right)^{5}\frac{r_{e}}{c}\,\omega^{2}.$ (82)
The value of ${\cal R}$ depends on the choice of a reference frame, which can be parameterized by the upstream motion $\beta_{\rm u}$ or $\kappa_{\rm u}$. Note that $\kappa_{\rm u}<1$ ($\beta_{\rm u}<0$) in frames moving in the $+z$ direction relative to the upstream plasma. The wave frequency $\omega$ in a chosen frame scales as $\omega\propto\kappa_{\rm u}$, so ${\cal R}\propto\kappa_{\rm u}^{2}$. Note also that the packet width $T$ depends on the choice of a frame as $\kappa_{\rm u}^{-1}$ (since $\omega T=inv$), and the radiative parameter $q\propto T{\cal R}/\kappa_{\rm u}$ is invariant.
@@
(1) Radiative losses $\dot{\gamma}_{e}$ may be important only when $w\gg 1$ because of the small value of ${\cal R}$ for real FRBs. We avoid simulating huge $w$ and $\kappa$ by rescaling ${\cal R}$ to larger values, so that we can experiment with the radiative regime $q\gg 1$ at moderate $w$ and $\kappa$ accessible to simulations. In sample models, we use the wave packet described by Equation (5) with $a_{\max}=(4/3)\sqrt{7}$ (which gives $w_{\max}=4$) and set the invariant radiative parameter $T{\cal R}/\kappa_{\rm u}=1$. This choice gives $q\sim 3$.
@@
(2) Both $\kappa$ and $\rho$ grow by a large factor during relaxation toward the steady state. The simulation is more efficient if performed in a boosted frame where $\kappa_{\rm u}\ll 1$. This helps avoid ultra-relativistic motions of the accelerated fluid along the wave propagation direction (large $\kappa\gg 1$) that would slow down the simulation. We use the frame where $\kappa_{\rm u}=0.05$. When presenting the results, we transform the fluid parameters to the upstream rest frame (where $\kappa_{\rm u}=1$), so figures will show $\kappa$ and $C=\rho/\rho_{\rm u}$ measured in the upstream frame.
Figure 6: Evolution of the compression peak $C_{\max}$ from the non-radiative steady state ($C_{\max}=C_{\max}^{\star}=13$ at $t=0$) after switching on radiative losses with $T{\cal R}/\kappa_{\rm u}=1$ (the radiative parameter ${\cal R}$ is defined in Equation (82)). The front evolves toward the new steady state with $C_{\max}^{\star}\approx 147$. The evolution is shown for three simulations with upstream magnetizations $\sigma_{\rm u}=3$, 10, 20. The initial and final states are independent of $\sigma_{\rm u}$ and shown in Figure 7.
@@
Figure 7: The initial and final profiles of the compression front in the three simulations shown in Figure 6. The initial state is the steady non-radiative front (Figure 3, red curve); here it is shown by the black dashed curve. Radiative losses make the front evolve until it reaches a new steady state (solid curve), which reproduces the solution in Equation (58).
@@
As an initial condition at $t=0$, we take the MHD fluid in the non-radiative steady front. In the rest frame of the upstream plasma, it is given by $\kappa_{0}(\xi)=1+w(\xi)$ and $C_{0}(\xi)=(\kappa_{0}^{2}+1)/2$. In particular, the initial compression peak is $C_{\max}(0)=13$. In the presence of radiative losses, this flow pattern is no longer a steady state, so it immediately begins to evolve at $t>0$. Both $\kappa$ and $C$ grow and eventually relax to a new steady state.
Figure 6 shows the evolution of $C_{\max}$, and Figure 7 shows the final profiles of $\kappa(\xi)$ and $C(\xi)$. The relaxation is determined by the chosen radiative coefficient ${\cal R}$ and the corresponding $q\sim 3$. In particular, $\kappa_{\max}$ grows from its initial value $51+w_{\max}=$ 5 to the final value $\kappa_{\max}=17$, consistent with the factor of $q\sim 3$, and $C_{\max}$ grows from 13 to $147$, confirming the expected increase $\sim q^{2}$.
@@
The simulations also confirm the expected scaling of the relaxation timescale $t_{\star}\propto q\sigma_{\rm u}$ at $q\gg 1$ and $\sigma_{\rm u}\gg 1$ (Equation 81). This scaling can be seen in Figure 6, which shows three models with $\sigma_{\rm u}=3$, 10, and 20. Note that $t_{\rm cross}\propto\kappa^{2}\propto q^{2}w^{2}$, and hence $t_{\star}$ should scale as $q^{3}$ at $q\gg 1$. One can see that the relaxation timescale is indeed increased by a large factor $\sim q^{3}\sim 30$ compared with the non-radiative front evolution shown in Figure 4.
@@
As expected, the final steady state is independent of $\sigma_{\rm u}$: it is the same in all three simulations with $\sigma_{\rm u}=3,10,20$. The steady-state profiles of $\kappa(\xi)$ and $C(\xi)$ are shown in Figure 7 and agree with Equation (58). Note that in the non-radiative regime ($q\ll 1$) the fluid comes back to rest as it exits behind the wave packet. In contrast, the fluid behind the radiative front has a large speed $\beta_{\rm D}>0$ ($\kappa\gg 1$). This is a consequence of momentum deposition by scattering during the fluid motion through the wave packet.
7 Parameters of the wind
@@
The obtained picture of a radio wave interaction with MHD fluid can now be applied to FRBs propagating through a radial wind. The wind serves as the upstream medium for the wave packet, and we begin with a brief summary of its parameters. They are denoted with subscript “u” for consistency of notation in this paper.
@@
The wind forms at the light cylinder $R_{\rm LC}=c/\Omega$ and its Lorentz factor grows as $\gamma_{\rm u}(r)\approx r/R_{\rm LC}$ until it reaches $\eta^{1/3}$; at larger radii, $\gamma_{\rm u}$ grows logarithmically (Lyubarsky & Kirk, 2001). We will use a simple approximation:
$\gamma_{\rm u}\approx\left\{\begin{array}[]{cc}\vspace*{2mm}\displaystyle{\frac{r}{R_{\rm LC}}}&\quad r<\eta^{1/3}R_{\rm LC}\\ \displaystyle{\left[\frac{\eta}{\ln 2}\ln\left(1+\frac{r}{\eta^{1/3}R_{\rm LC}}\right)\right]^{1/3}}&\quad r>\eta^{1/3}R_{\rm LC}\\ \end{array}\right.$ (85)
@@
Note that $L_{\rm w}$ determines the wind magnetic field in the static lab frame $B_{\rm u}(r)$ according to Equation (83). The corresponding magnetic field in the rest frame of the wind (ahead of the FRB wave packet) is $\tilde{B}=B_{\rm u}/\gamma_{\rm u}$. It sets another important parameter — the gyro-frequency of the wind plasma,
@@
$\alpha\equiv\frac{\tilde{\omega}_{B}^{\rm u}}{\tilde{\omega}_{\rm u}}\approx\frac{2e}{mcr\omega}\left(\frac{L_{\rm w}}{c}\right)^{1/2}\approx 0.1\,\frac{L_{\rm w,37}^{1/2}}{\nu_{9}r_{12}}.$ (88)
@@
Below we examine the compression front driven by an FRB emitted at some radius $r$ and propagating outward through the wind. We first describe the front at large $r$, where FRBs induce heating in the regular regime ($\tilde{\omega}_{B}<\tilde{\omega}$), and then consider FRBs emitted at smaller $r$, where heating is stochastic and more dramatic effects occur, including damping of the FRB.
8 Regular heating zone
@@
Strong compression is expected at radii where the FRB has $a>1$. The wave strength parameter $a$ (Equation 1) may be written as $a=r_{1}/r$, where
$r_{1}=\left(\frac{r_{e}L}{mc\,\omega^{2}}\right)^{1/2}\approx 1.6\times 10^{13}\,L_{42}^{1/2}\nu_{9}^{-1}\,{\rm cm}.$ (89)
8.1 Saturated compression ($r>r_{\star}$)
@@
One can quickly formulate a compression front model using the steady-state solution described by Equations (52), (57), and (75), assuming the regular heating regime ($\tilde{\omega}_{B}<\tilde{\omega}$) and neglecting radiative losses. Using the definition of the proper compression factor $\tilde{C}\equiv\tilde{\rho}/\tilde{\rho}_{\rm u}$, the steady front solution can be stated as
$\tilde{C}_{\star}=\frac{\kappa_{\star}}{\kappa_{\rm u}}=1+\frac{3}{2}\left(\sqrt{a^{2}+1}-1\right).$ (90)
@@
The corresponding compression factor for the wind density measured in the lab frame, $C\equiv\rho/\rho_{\rm u}$, is
$C_{\star}=\frac{\kappa_{\star}^{2}+1}{\kappa_{\rm u}^{2}+1}\approx\tilde{C}_{\star}^{2},$ (91)
@@
where $\kappa\equiv\gamma_{\rm D}(1+\beta_{\rm D})$ describes the fluid drift speed $\beta_{\rm D}$, and we used $\kappa_{\rm u}\gg 1$ since the upstream medium is a relativistic wind. Outside the radius $r_{1}$ (where $a\ll 1$), one finds weak compression $C_{\star}\approx 1$ while inside $r_{1}$ (where $a\gg 1$) one finds $C_{\star}\approx(9/4)a^{2}=(9/8)a_{0}^{2}\approx a_{0}^{2}$. A simple approximate expression matching the two cases is
$C_{\star}\approx 1+a_{0}^{2}.$ (92)
Remarkably, this compression is consistent with the results of Sobacchi et al. (2024), who investigated a steady plasma flow in a (nearly monochromatic) wave packet using a different technique.
@@
However, the simple solution stated in Equation (92) does not correctly describe compression fronts in FRBs at small radii $r\ll r_{1}$, for two reasons:
@@
(1) At $r<r_{\star}$ defined below, the compression front does not reach the steady state, so the actual compression factor is well below its saturated value, $C\ll C_{\star}$.
@@
(2) Compression promotes the transition to stochastic acceleration of plasma particles in the FRB wave packet. The plasma then experiences much stronger heating, and the momentum exchange between the FRB wave packet and the plasma is enhanced, changing $\kappa$ and $C$.
These two features qualitatively change the picture of the FRB compression front, as described below.
8.2 Unsaturated compression $(r<r_{\star})$

### v2

@@ VI.2 Evolution of a radiative compression front
If $q\gg 1$, the fluid element moving through the wave packet experiences strong radiative losses. In this case, the radio wave continually deposits energy to offset the losses and keep the average $\tilde{\gamma}=\sqrt{1+a^{2}}$ (B22). The deposited power per particle in the fluid frame approximately equals the average emitted power $\delta\tilde{{\cal E}}_{e}/\delta\tilde{t}$ (given by Equation (D9) with $\beta_{\rm D}=0$), so the wave-fluid interaction at $q\gg 1$ is essentially a scattering process.
The scattering exerts a force on the fluid. The wave passes momentum with rate $\delta\tilde{{\cal E}}_{e}/c\delta\tilde{t}$, and the fluid radiates momentum with rate $\delta\tilde{{\cal P}}_{e}^{z}/\delta\tilde{t}$, so the net rate of momentum deposition into the MHD fluid is
$\frac{\delta\tilde{{\cal P}}}{\delta\tilde{t}}=\frac{\,\delta\tilde{{\cal E}}_{e}}{c\delta\tilde{t}}-\frac{\delta\tilde{{\cal P}}_{e}^{z}}{\delta\tilde{t}}$. (74)
Transformation from the local fluid frame to a lab frame gives $\delta{\cal P}/\delta t=\delta\tilde{{\cal P}}/\delta\tilde{t}$ since $\delta t=\gamma\delta\tilde{t}$ and $\delta{\cal P}=\gamma\delta\tilde{\cal P}$ (from Lorentz transformation, using $\delta\tilde{\cal E}/\delta\tilde{\cal P}=\tilde{v}_{\rm D}=0$). Thus, Equation (74) gives the Lorentz-invariant force $\dot{\cal P}_{\rm sc}=\delta{\cal P}/\delta t$ applied to the MHD fluid by the scattering of the radio wave. Substitution of Equations (D9) and (D10) (with $\beta_{\rm D}=0$ in frame $\tilde{\cal K}$) yields
$\dot{\cal P}_{\rm sc}=\frac{\sigma_{\rm T}U}{\kappa^{2}}(1+a^{2}).$ (75)
@@
with $a\sim a_{\max}$. This sets the budget for the fluid momentum gain per crossing time $t_{\rm cross}$. The gained momentum is shared with the magnetic field frozen in the fluid, increasing ${\cal P}_{\rm MHD}$.
@@
where $q$ may be evaluated using the column energy $\Sigma$ in the middle of the wave packet (Equation 57). Equation (77) gives the timescale for approaching the steady state, $t_{\star}=t_{\cal P}(\kappa_{\star})$ if we substitute $\kappa\sim\kappa_{\star}=(1+q)(1+w)$ (Equation 56) and $h_{\star}=1+w+\kappa_{\star}\sigma_{\rm u}\approx\kappa_{\star}\sigma_{\rm u}$. Interpolating between the limits of $q\ll 1$ and $q\gg 1$, we find
$t_{\star}\sim(1+q)\,\sigma_{\rm u}t_{\rm cross}^{\star}\qquad(\sigma_{\rm u}\gg 1).$ (78)
@@
Front relaxation to the steady state in the radiative regime is found by solving Equations (46) and (47) similarly to section VI.1, but now keeping the term $n\dot{\cal Q}^{-}_{e}/c^{2}$. Using Equation (50) and $U=m\,\omega^{2}a^{2}/4\pi r_{e}$ (Equation 28), we rewrite $\dot{\cal Q}^{-}_{e}$ as follows
$\frac{\dot{\cal Q}^{-}_{e}}{mc^{2}}={\cal R}\,\frac{a^{2}(1+a^{2})}{\gamma_{\rm D}\kappa^{3}},\qquad{\cal R}\equiv\frac{2r_{e}}{3c}\,\omega^{2}.$ (79)
[…]
$q(T)=\frac{{\cal R}}{\kappa_{\rm u}}\int_{0}^{T}a^{2}d\xi=\frac{3}{8}a_{\max}^{2}\frac{T{\cal R}}{\kappa_{\rm u}}.$ (80)
A characteristic value of $q$ in the middle of the front is $q(T/2)=q(T)/2$.
[…]
@@
[…]
@@
(2) Both $\kappa$ and $\rho$ grow by a large factor during relaxation toward the steady state. The simulation more efficiently performed in a boosted frame where $\kappa_{\rm u}\ll 1$. This helps avoid ultra-relativistic motions of the accelerated fluid along the wave propagation direction (large $\kappa\gg 1$) that would slow down the simulation. We use the frame where $\kappa_{\rm u}=0.05$. When presenting the results, we transform the fluid parameters to the upstream rest frame (where $\kappa_{\rm u}=1$), so figures will show $\kappa$ and $C=\rho/\rho_{\rm u}$ measured in the upstream frame.
[…]
@@
Figure 8: Initial (dashed) and final (solid) profiles of the compression front in the three simulations shown in Figure 7. The initial state is the steady non-radiative front (Figure 4, red curve). Radiative losses make the front evolve until it reaches the new steady state, which reproduces the analytical solution (Equation 56).
@@
As an initial condition at $t=0$, we take the steady flow pattern in the non-radiative front. In the rest frame of the upstream plasma, it is given by $\kappa_{0}(\xi)=\sqrt{1+a^{2}(\xi)}$ and $C_{0}(\xi)=1+a^{2}(\xi)/2$, so the initial compression peak is $C_{\max}(0)=13$. In the presence of radiative losses, this flow pattern is no longer a steady state, and it immediately begins to evolve at $t>0$. Both $\kappa$ and $C$ grow and eventually relax to a new steady state.
[…]
@@
The simulations also confirm the expected scaling of the relaxation timescale $t_{\star}\sim(1+q)\sigma_{\rm u}t_{\rm cross}^{\star}$ at $\sigma_{\rm u}\gg 1$ (Equation 78). The scaling $t_{\star}\propto\sigma_{\rm u}$ is seen in Figure 7, which shows three models with $\sigma_{\rm u}=3$, 10, and 20. Note also that $t_{\rm cross}^{\star}\propto\kappa_{\star}^{2}\propto(1+q)^{2}$ and hence $t_{\star}$ is expected to scale as $(1+q)^{3}$. We observe that the relaxation timescale is indeed increased by a large factor of $\approx(1+q)^{3}/2$ compared to the non-radiative front relaxation shown in Figure 6.
@@
As expected, the final steady state (Figure 8) is independent of $\sigma_{\rm u}$: it is the same in all three simulations with $\sigma_{\rm u}=3,10,20$. The observed final state satisfies Equation (56) with $\kappa_{\rm u}=1$ (since we view the results in the upstream rest frame). Note that in the non-radiative regime ($q\ll 1$) fluid comes back to rest as it exits behind the wave packet. By contrast, fluid behind the radiative front ($q\gg 1$) retains a large speed $\beta_{\rm D}\sim 1$ ($\kappa\gg 1$). This is the result of irreversible momentum deposition by the scattering of the radio wave.
VII Parameters of the wind
@@
The obtained picture of the radio wave interaction with MHD fluid can now be applied to FRBs propagating through a radial wind carrying a transverse magnetic field, with some magnetization parameter $\sigma_{\rm u}$. The wind serves as the upstream medium for the radio wave packet, and we begin with a brief summary of its parameters. They are denoted with subscript “u” for consistency of notation in this paper.
@@
The wind forms at the light cylinder $R_{\rm LC}=c/\Omega$ and its Lorentz factor grows as $\gamma_{\rm u}(r)\approx r/R_{\rm LC}$ until it reaches $\eta^{1/3}$; at larger radii, $\gamma_{\rm u}$ grows logarithmically (Beskin et al., 1998; Lyubarsky & Kirk, 2001). We will use a simple approximation:
[…]
@@
Note that $L_{\rm w}$ determines the wind magnetic field in the static lab frame $B_{\rm u}(r)$ according to Equation (81). The corresponding magnetic field in the rest frame of the wind (ahead of the FRB wave packet) is $\tilde{B}_{\rm u}=B_{\rm u}/\gamma_{\rm u}$. It sets another important parameter — the gyro-frequency of the wind plasma,
@@
$b_{\rm u}\equiv\frac{\tilde{\omega}_{B}^{\rm u}}{\tilde{\omega}_{\rm u}}=\frac{r_{b}}{r},\quad\;\,r_{b}=\frac{2eL_{\rm w}^{1/2}}{mc^{3/2}\omega}\approx 10^{11}\frac{L_{\rm w,37}^{1/2}}{\nu_{9}}\,{\rm cm}.$ (86)
@@
Below we examine the compression front driven by an FRB emitted at some radius $r$ and propagating outward through the wind. We first describe the front at large $r$, where plasma responds to the radio wave in the regular oscillation regime. Then, we consider FRBs emitted at smaller $r$, where stochastic heating occurs and leads to more dramatic effects, including damping of the FRB.
VIII Regular oscillation zone
@@
Strong compression is expected at radii where the FRB wave packet has $a_{\max}>1$. The strength parameter $a_{\max}$ (Equation 1) may be written as
$a_{\max}=\frac{r_{1}}{r},$ (87)
where
$r_{1}=\left(\frac{r_{e}L}{mc\,\omega^{2}}\right)^{1/2}\approx 1.6\times 10^{13}\,L_{42}^{1/2}\nu_{9}^{-1}\,{\rm cm},$ (88)
and $L$ is the peak luminosity of the FRB.
VIII.1 Saturated compression ($r>r_{\star}$)
@@
One can quickly formulate a compression front model using the steady-state solution described by Equations (51) and (58), which assumes regular oscillations of the plasma in the radio wave. The shape of the wave packet $a(\xi)$ then determines the profile of the proper compression factor in the front, $\tilde{C}\equiv\tilde{\rho}/\tilde{\rho}_{\rm u}$,
$\tilde{C}_{\star}=\frac{\kappa_{\star}}{\kappa_{\rm u}}=\sqrt{1+a^{2}},$ (89)
@@
where star indicates that the solution is stated for the steady front. The corresponding compression of the wind density measured in the lab frame, $C\equiv\rho/\rho_{\rm u}$, is
$C_{\star}(\xi)=\frac{\kappa_{\star}^{2}+1}{\kappa_{\rm u}^{2}+1}\approx 1+a^{2},$ (90)
@@
where $\kappa\equiv\gamma_{\rm D}(1+\beta_{\rm D})$ describes the fluid drift speed $\beta_{\rm D}$, and we used $\kappa_{\rm u}\gg 1$ since the upstream medium is a relativistic wind. The maximum compression in the steady-state front is at the peak of the wave packet: $C_{\star}^{\max}\approx 1+a_{\max}^{2}$.
@@
The simple solution stated in Equation (90) is valid only at sufficiently large radii $r>r_{\star}$ evaluated below. It does not describe compression fronts for FRBs at smaller radii for two reasons:
@@
(1) At $r<r_{\star}$, the compression front does not reach the steady state (section VIII.2). Then, the actual compression factor is well below its saturated value, $C\ll C_{\star}$.
@@
(2) Compression promotes the transition to stochastic heating of plasma particles by the radio wave (section IX). The plasma then receives much more energy (and hence much more momentum), which changes $\kappa$ and $C$.
VIII.2 Unsaturated compression $(r<r_{\star})$